solve the equation by a method of your choice.
1. x^2 -2x -15=0

Write a quadratic equation that has the given solutions.
1. -2 and 5
2. -3 and -5

Answer by stanbon(75887)

solve the equation by a method of your choice.
1. x^2 -2x -15=0
(x-5)(x+3) = 0
x = 5 or x = -3
============================
Write a quadratic equation that has the given solutions.
1. -2 and 5
y = (x+2)(x-5)
y = x^2-3x-10
--------------------
2. -3 and -5
y = (x+3)(x+5)
y = x^2+8x+15
=============================
Cheers,
Stan H.
================